OpenStudy (anonymous):

Which expression is equivalent to2^−5? A. 1/5×1/5 (Fraction) B. (−5)(−5) C. 1/2×1/2×1/2×1/2×1/2 (fraction) D. (−2)(−2)(−2)(−2)(−2)

OpenStudy (anonymous):

@YesThisIsDog

OpenStudy (caominhim):

2-5 = 3 none of those equal 3

OpenStudy (anonymous):

I think its D @YesThisIsDog

OpenStudy (caominhim):

-2^5 = -32

OpenStudy (anonymous):

That is correct.

OpenStudy (caominhim):

wait no \[2^{-5} = \frac{ 1 }{ 2^{5} }\]

OpenStudy (caominhim):

its C

OpenStudy (anonymous):

Are you sure?

OpenStudy (caominhim):

yea, if you apply a negative power it flips it\[x^{-n} = \frac{ 1 }{ x^{n} }\]